OpenStudy (anonymous):

A company sells cardboard that is nine and one-eighth millimeters thick. Write an equation to solve for the number of sheets of cardboard, c, in a pile that is two hundred fifty-five and one-half millimeters thick. What must be done to both sides to solve the equation?

OpenStudy (minecrafter64):

you there?

OpenStudy (texaschic101):

(9 1/8)c = 255 1/2 -- change to improper fractions (73/8)c = 511/2 -- divide both sides by 73/8 c = (511/2) / (73/8) c = 511/2 * 8/73 c = 4088/146 c = 28 sheets